OpenStudy (anonymous):

please answer my question ,i need the help its very difficult,(4^3/5.2^4/5)^5/12

OpenStudy (anonymous):

gosh!!,i love math,sometimes i hate math,,coz there's a time comes that math is very difficult

OpenStudy (anonymous):

you meant \[(4^{\frac{3}{5}}\times 2^{\frac{4}{5}})^{\frac{5}{12}}\] right?

OpenStudy (anonymous):

so again gimmick is to write \[4=2^2\] so \[4^{\frac{3}{5}}=(2^2)^{\frac{3}{5}}=2^{\frac{6}{5}}\]

OpenStudy (anonymous):

so you have \[2^{\frac{6}{5}}\times 2^{\frac{4}{5}}=2^{\frac{10}{5}}=2^2=4\]

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

yw but hold on let me check my work because it is early

OpenStudy (anonymous):

ukie

OpenStudy (anonymous):

yeah it is right but you are not done because you still have to raise to the power of 12/5 so you get \[(2^2)^{\frac{5}{12}}=2^{\frac{5}{6}}=\sqrt[6]{2^5}\]

OpenStudy (anonymous):

thank you,.can i ask 1 more problem??

OpenStudy (anonymous):

ask away

OpenStudy (anonymous):

\[(2x^3/5)^5\div (5y^4/5\])^5

OpenStudy (anonymous):

\[(\frac{2x^3}{5})^5\times (\frac{5}{5y^4})^5\] is a start

OpenStudy (anonymous):

"multiply by the reciprocal"

OpenStudy (anonymous):

which i should mutiply by reciprocal???the last fraction??

OpenStudy (anonymous):

yes i flipped it right?

OpenStudy (anonymous):

now you can combine as one term to the power 5 if you like (since they are both raised to the fifth power) and get \[(\frac{10x^3}{25y^4})^5\]

OpenStudy (anonymous):

then maybe cancel a 5 inside the parentheses ( i should have done this first) to get \[(\frac{2x^3}{5y^4})^5\]

OpenStudy (anonymous):

i would leave it like that but you could also write \[\frac{2^5x^1{5}}{5^5y^{20}}\] if you like

OpenStudy (anonymous):

ukie thank you!!!

OpenStudy (anonymous):

yw